OpenStudy (anonymous):

What is the y-intercept of y = 2(3)^x?

OpenStudy (anonymous):

substitute 0 to x if you want to solve for y-intercept.

OpenStudy (unklerhaukus):

\[y(x) = 2(3)^x\]\[y(0) = 2(3)^0=?\]

OpenStudy (anonymous):

(0, 2)

OpenStudy (anonymous):

any number powered to 0 is 1

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

yep (0,2) is the y-intercept .